### v2

# TCM-FTP: Fine-Tuning Large Language Models for Herbal Prescription Prediction

Abstract
Traditional Chinese medicine (TCM) has relied on specific combinations of herbs in prescriptions to treat various symptoms and signs for thousands of years. Predicting TCM prescriptions poses a fascinating technical challenge with significant practical implications. However, this task faces limitations due to the scarcity of high-quality clinical datasets and the complex relationship between symptoms and herbs. To address these issues, we introduce DigestDS, a novel dataset comprising practical medical records from experienced experts in digestive system diseases. We also propose a method, TCM-FTP (TCM Fine-Tuning Pre-trained), to leverage pre-trained large language models (LLMs) via supervised fine-tuning on DigestDS. Additionally, we enhance computational efficiency using a low-rank adaptation technique. Moreover, TCM-FTP incorporates data augmentation by permuting herbs within prescriptions, exploiting their order-agnostic nature. Impressively, TCM-FTP achieves an F1-score of 0.8031, significantly outperforming previous methods. Furthermore, it demonstrates remarkable accuracy in dosage prediction, achieving a normalized mean square error of 0.0604. In contrast, LLMs without fine-tuning exhibit poor performance. Although LLMs have demonstrated wide-ranging capabilities, our work underscores the necessity of fine-tuning for TCM prescription prediction and presents an effective way to accomplish this.
Index Terms:
Large language models, Traditional Chinese medicine, Fine-tuning, Prescription prediction, Herb dosage prediction
I INTRODUCTION
Traditional Chinese Medicine (TCM), integral to Chinese healthcare for millennia, encompasses practices such as herbal medicine, acupuncture, cupping therapy, and tuina massage. As the primary modality of TCM, herbal medicine has demonstrated effectiveness in treating COVID-19, leading to improved cure rates and reduced mortality [1]. Herbal prescriptions in TCM require doctors to assess patient symptoms using four diagnostic methods—observation (wang), listening and smelling (wen), questioning (wen), and pulse-taking (qie)—guided by the principle of li-fa-fang-yao [2]. The extensive experience and continuous practical feedback required for training result in prolonged training periods, contributing to a shortage of practitioners and limited TCM resources. Developing an effective prescription predictor could alleviate challenges such as doctor shortages and lengthy training requirements.
Computational systems for prescription prediction in TCM model the complex relationships between symptoms and herbs [3]. Language generation techniques, treating prescription generation as a machine translation problem using sequence-to-sequence (seq2seq) models [4], have shown promise. Despite positive results from generative models [5, 6, 7, 8], current models still face the following limitations: (1) Scarcity of clinical datasets. Current models often rely on datasets derived from classical documents instead of high-quality clinical records, introducing noise and limiting the effectiveness of personalized prescription recommendations [9]. Moreover, there is a significant lack of high-quality clinical data for prescription prediction [10]. (2) Sub-optimal Performance. Existing models often underperform and may mislead users due to low prediction accuracy [10]. Relying solely on clinical data is insufficient to enhance model performance [11]; integrating advanced language modeling techniques is necessary to improve the representation and encoding capabilities of prescription prediction models. (3) Lack of herb dosage. Existing models overlook dosage weights, a critical component in TCM for effective treatment [3]. Furthermore, there is a lack of evaluation metrics specifically designed for herb dosage prediction.
Motivated by the robust predictive capabilities of large language models (LLMs), we develop a high-quality prescription dataset, DigestDS, derived from clinical medical records, and propose TCM-FTP —a novel LLM-based TCM prescription generation model (Fig. 1). Specifically, we employ a low-rank adaptation technique (LoRA) [12] for efficient fine-tuning of the LLM. Exploiting the order-agnostic nature of herb prescriptions, we implement data augmentation by randomizing the sequence of herbs in the training data. Validated on DigestDS, our model achieves a precision of 0.7951, a recall of 0.8113, and an F1-score of 0.8031, significantly outperforming previous methods. Our main contributions are outlined as follows:
Figure 1: Workflow of the TCM-FTP. Our work consists of four parts: “Data Collection” involves gathering and organizing raw data; “Data Processing” includes data preprocessing, prompt design, and integrating data augmentation to create a fine-tuning dataset; “Fine-tuning” utilizes the ShenNong LLM and LoRA technique to optimize the model; and “Evaluation” assesses the outcomes using both quantitative and qualitative evaluation metrics.
  • We construct DigestDS, a TCM clinical diagnostic dataset derived from real-world clinical records, serving as a foundational resource for understanding the nuanced relationships between symptoms and effective prescriptions in TCM.
  • We introduce TCM-FTP, a novel LLM-driven method for generating TCM prescriptions that excels at capturing the intricate interplay between symptoms and herbal treatments, marking a significant advancement in prescription prediction. Additionally, we introduce the Normalized Mean Square Error (NMSE) as an evaluation metric for dosage prediction, pioneering the first dosage prediction metric tailored to real-world scenarios in this field.
  • We empirically validate TCM-FTP against existing models on our dataset, demonstrating its superior performance in predicting herbs and dosages. Expert case studies by TCM professionals further provide valuable insights into the adaptability and robustness of our innovative approach.
II BACKGROUND AND SIGNIFICANCE
II-A Problem Definition
We now formally define the prescription prediction problem. Given a prescription dataset $\mathcal{P}_{\text{train}}$, an element in $\mathcal{P}_{\text{train}}$ consists of a symptom description $s$ and its corresponding prescription $\{h_{i},w_{i}\}^{i\in[k]}$. Here, $h$ denotes the herb name, $w$ represents the herb dosage, and $[k]$ signifies a list ranging from 1 to $k$. Our objective is to train a model $\mathcal{M}$ such that $\mathcal{M}(s)$ reproduce $\{h_{i},w_{i}\}^{i\in[k]}$ accurately. For alignment with the language generation task, we concatenate $\{h_{i},w_{i}\}^{i\in[k]}$ using a comma separator to form a single sentence. Concrete examples are presented in Fig. 1.
II-B TCM Herbal Prescription Prediction
Research on TCM prescription prediction mainly falls into three categories: topic model-based, graph model-based, and language model-based approaches. Topic model-based approaches treat relationships between symptoms and herbs as that of documents and topics [13]. These approaches rely on statistical relationships, lacking a semantic understanding of symptoms. Graph-based approaches construct a medical knowledge graph to model the relationships between symptoms and herbs [14]. However, these approaches also lack consecutive semantic information regarding symptoms. Language model-based prescription prediction models take patient symptom descriptions as input and generate herbal prescriptions sequentially. TCM Translator [15] uses transformer architectures to distill context vectors from symptoms and LSTM as the decoder. AttentiveHerb [5] employs a seq2seq model with dual attention mechanisms to distinguish primary from secondary symptoms and map herb-symptom interactions using clinical data. Herb-Know [6] utilizes herb descriptions to model associations with symptoms and evaluates whether herb effects align with symptom descriptions, and other related work like TCMBERT [7], RoKEPG [8] also have shown their effectiveness on TCM prescription prediction.
II-C Large Language Models in TCM
LLMs have advanced significantly in NLP tasks, with models like ChatGPT and GPT-4 [16] drawing attention, though OpenAI has not disclosed architectural details. Open-source LLMs, such as LLaMA [17] and ChatGLM [18], have spurred research interest, though they often struggle in specialized fields like medicine or finance. Supervised fine-tuning has become the standard to enhance these models with domain-specific knowledge. In TCM, tailored models have emerged. Bentsao [19] fine-tunes LLaMA using structured and unstructured knowledge from CMeKG [20]. Huatuo [21] incorporates ChatGPT data with Reinforcement Learning from Artificial Intelligence Feedback (RLAIF) and Proximal Policy Optimization (PPO). Zhongjing [22] applies continual pretraining with supervised fine-tuning on a Chinese medical dialogue dataset, integrating Reinforcement Learning from Human Feedback (RLHF). ShenNong [23] combines LoRA-based fine-tuning on instructional data and TCM knowledge, supported by over 110,000 instructions.
II-D Parameter Efficient Fine-Tuning
Parameter Efficient Fine-Tuning (PEFT) utilizes a small amount of parameters to fine-tune a large language model effectively. Assuming there is a pretrained model $f_{\theta}(y|x)$, PEFT seeks to adjust a limited number of parameters, $\Delta\theta$, such that $|\theta|\gg|\Delta\theta|$. In contrast to conventional fine-tuning, which updates all parameters, denoted as $|\theta|=|\Delta\theta|$, and requires significant computational resources, PEFT selectively updates a limited number of learnable parameters to achieve results comparable to complete fine-tuning. In this work, we employ LoRA [12] to efficiently fine-tune LLMs on the target dataset.
III MATERIALS AND METHODS
In this study, as illustrated in Fig. 1, we gather a high-quality prescription dataset DigestDS and propose a PEFT approach for prescription prediction TCM-FTP. DigestDS is comprised of practical medical records collected from specialists in digestive system disorders. TCM-FTP utilizes supervised fine-tuning to train the model on DigestDS, incorporating LoRA technique and an effective data augmentation technique, which involves permuting the herbs in the prescriptions.
III-A Datasets
III-A1 Data collection
We collect outpatient medical record data from practical patient records generated by specialists in TCM hospital 111by Dr. Runshun Zhang in Guang’anmen Hospital, China over a span from 2008 to 2022. The prescriptions specifically focus on digestion system disorders.
III-A2 Data processing
Initially, we remove the incomplete data items and erase any privacy information. Subsequently, we exclude irrelevant information, retaining only essential information for prescription prediction. Specifically, we keep the chief complaint, medical history, and tongue-coating details from the symptom descriptions, as well as the names and dosages of herbs in the prescriptions.
III-A3 Data statistics
We present the processed data statistics in Table I. Prescription items are randomly divided into training (90%) and testing (10%) datasets. The training dataset comprises 16,896 samples with an average of 23.92 herbs per prescription, while the test dataset includes 2,057 samples averaging 23.86 herbs per prescription. Fig. 2 illustrates the distributions of herb counts in the training and testing datasets.
TABLE I: Statistics of training and test datasets The table displays the median, mean, and standard deviation for the number of herbs per prescription. Category: Number of distinct herbs in the dataset.
dataset size category median mean std
training 16,896 674 24 23.92 5.69
test 2,057 533 24 23.86 5.44
(a) Training dataset
(b) Test dataset
Figure 2: Distribution of the number of herbs in prescriptions The blue curves represent kernel density estimates.
III-B TCM-FTP
To better model the intricate relationships between symptoms and herbs, we propose TCM-FTP, which employs a pre-trained LLM coupled with an auto-regressive mechanism for TCM prescription prediction. For efficient fine-tuning of the LLM, we employ the LoRA technique[12], which optimizes the LLM with limited parameters and computational resources. To enhance the dataset, we permute the herbs in prescriptions to leverage the order-agnostic property inherent in TCM prescriptions.
Given a prescription dataset $\mathcal{P}_{train}$ with $\{s,\{h_{i},w_{i}\}^{i\in[k]}\}$, our primary goal is to use a language model to minimize the negative logarithm of the estimation of conditional probability:
$\min\limits_{\theta}\mathop{\mathbb{E}}\limits_{s,\{h_{i},w_{i}\}^{i\in[k]}\sim\mathcal{P}_{train}}\sum\limits_{i}^{k}-\log P((h_{i},w_{i})|h_{<i},w_{<i};s;\theta),$ (1)
where $s$ is the symptom description, $k$ is the number of herbs in the prescription, and $(h_{i},w_{i})$ represents the i-th (herb, dosage) pair with $i\in\{1,2,...,k\}$. $h_{<i}$ and $w_{<i}$ represent a set of herbs and a set of weights with index less than $i$, respectively. $\theta$ represents our model parameters.
We concatenate the herbs and dosages into one string to use the auto-regressive mechanism in LLMs and get the goal:
$\min\limits_{\theta}\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}s,\{h_{i},w_{i}\}^{i\in[k]}\sim\mathcal{P}_{train}\\ y=\textsc{cat}(\{h_{i},w_{i}\}^{i\in[k]})\end{subarray}}\sum\limits_{j=1}^{|y|}-\log P(y_{j}|\{y_{<j}\};s;\theta),$ (2)
where $\{y_{<j}\}$ denotes the set of the tokens from index $1$ to index $j-1$ in $y$, and $|y|$ represents the number of tokens in $y$. $\textsc{cat}(\{h_{i},w_{i}\}^{i\in[k]})$ is the concatenated string.
III-B1 Low-rank Adaptation
We employ the LoRA [12] to reduce computational resources. Towards $f_{\theta}$, we encounter two primary types of parametric functions within our model framework: linear and embedding functions, defined as:
${\mathcal{F}_{\text{linear}}}(x)={W_{l}}\cdot x,{\mathcal{F}_{\text{emb}}}(x)=\text{EMB}\left(x;{W_{e}}\right),$ (3)
where EMB$(x;\cdot)$ represents the embedding operator that selects the $x$-th column from the specified matrix, and $W_{l}$ and $W_{e}$ denote the parameters for the linear and embedding functions, respectively. For fine-tuning, we introduce updates via low-rank matrices $A$ and $B$ to modify the original functions:
$\begin{split}{\mathcal{F}_{\text{linear}}}{(x)^{\prime}}&={\mathcal{F}_{\text{linear}}}(x)+{A_{l}}\cdot{B_{l}}\cdot x,\\ {\mathcal{F}_{\text{emb}}}{(x)^{\prime}}&={\mathcal{F}_{\text{emb}}}(x)+{A_{e}}\cdot\text{EMB}\left(x;{B_{e}}\right),\end{split}$ (4)
where $A_{l}$, $B_{l}$, $A_{e}$, and $B_{e}$ are the learnable parameters. With the updates to the low-rank matrices represented by $\Delta\theta$, our goal is:
$\min\limits_{\Delta\theta}\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}s,\{h_{i},w_{i}\}^{i\in[k]}\sim\mathcal{P}_{train}\\ y=\textsc{cat}(\{h_{i},w_{i}\}^{i\in[k]}\end{subarray}}\sum\limits_{j=1}^{|y|}-\log P(y_{j}|\{y_{<j}\};s;\theta+\Delta\theta).$ (5)
III-B2 Order-Agnostic Property
[…]
III-C Baselines
To demonstrate the performance of our model, we compare it with several baselines, including TCM prescription prediction models (PTM [3], TCMPR [24], KDHR [25], and PresRecST [10]), and pre-trained language models (Mengzi (T5-base) [26], GPT-3.5 and GPT-4.0 [16]).
III-D Experimental setup
We implement the fine-tuning process with the Transformers package from Hugging Face [^2] https://huggingface.co/docs/transformers. The number of training epochs is set to 10. We adopt a cosine scheduler, with the LoRA rank set to 16 and the LoRA alpha to 32. We consider the number of herb permutations, K, as 20 and 50, respectively. The default learning rate is set to $1\times 10^{-3}$. We leverage the gradient accumulation technique to increase the batch size, with 8 accumulation steps and a batch size of 16. We employ two foundation models Chinese-Alpaca-Plus-7B  [^3] https://github.com/ymcui/Chinese-LLaMA-Alpaca/tree/main and ShenNong  [^4] https://huggingface.co/michaelwzhu/ShenNong-TCM-LLM. Chinese-Alpaca-Plus-7B is a variant of LLaMA [17] with continual pre-training and supervised fine-tuning on Chinese corpus, denoted as LLaMA+ for simplicity. ShenNong is a further refinement of LLaMA+, fine-tuned with TCM instruction datasets. We use an 8-V100 GPU machine for fine-tuning and the running time for TCM-FTP ($K=50$) is 146 hours. During the inference stage, We employ top-k and top-p combinations for decoding, setting the top-k to 50 and the top-p to 0.7, with a default temperature of 0.95.
III-E Evaluation Metrics
Quantitative Evaluation.
  • Precision, recall, and F1-score: Precision is the proportion of correct herb predictions, recall is the proportion within true herbs, and F1-score is their harmonic mean.
  • Normalized mean square error (NMSE): We design a metric to evaluate the accuracy of predicted dosage by normalizing the squared differences using the original weights. For a given prescription $p$ and a predicted prescription $\hat{p}$, we suppose they are composed of a set of pairs of a herb and a dosage, $p=\{(h,w)\}$, where $h,w$ refer to herb and dosage respectively, and $\hat{p}=\{(h^{\prime},w^{\prime})\}$ indicates the generated result from model. The NMSE is defined as: $NMSE=\frac{1}{Z}\sum_{\begin{subarray}{c}(h,w)\in p,\\ \left(h^{\prime},w^{\prime}\right)\in\hat{p}\end{subarray}}\mathbf{1}\left[h=h^{\prime}\right]\left(\frac{w^{\prime}-w}{w}\right)^{2},$ where $Z$ is the number of correctly predicted herbs, and $\mathbf{1}[\cdot]$ is the indicator function. For baseline, we use the average dosage of each herb from the training data for dosage predictions.
Qualitative Evaluation. Existing quantitative evaluation metrics assess model quality based solely on sample labels, neglecting the compatibility and symptom-specific effectiveness of prescriptions. To comprehensively evaluate our model, we engaged five TCM experts to conduct an expert qualitative evaluation (EQE) of selected prescriptions generated by our model. Each doctor independently assessed the prescriptions for herbal effectiveness in treating symptoms (SHE) and herbal compatibility (HC), assigning scores on a scale of 0 to 5. Higher scores indicate greater effectiveness or compatibility.
IV RESULTS AND DISCUSSION
IV-A Overall Results
TABLE II: Prediction results of herbs and dosages in prescriptions TCM-FTP is evaluated with different foundation models.
Category Model Precision Recall F1-score NMSE NMSEbase
TCM prescription prediction models PTM [3] 0.5372 0.5777 0.5567 - -
TCMPR [24] 0.5241 0.4570 0.4882 - -
KDHR [25] 0.4917 0.3898 0.4349 - -
PresRecST [10] 0.5061 0.4016 0.4419 - -
[…]
GPT-4.0 0.0605 0.0761 0.0111 - -
Mengzi (T5-base) [26] 0.7332 0.7474 0.7403 0.0754 0.1378
[…]
TCM-FTP (llama+, K=50) 0.7916 0.8118 0.8016 0.0619 0.1462
TCM-FTP (ShenNong, K=20) 0.7919 0.8100 0.8008 0.0607 0.1441
TCM-FTP (ShenNong, K=50) 0.7951 0.8113 0.8031 0.0604 0.1431
As shown in Table II, the proposed TCM-FTP outperforms all baseline models in the herb prediction task on DigestDS, achieving an F1-score of 0.8016 using LLaMA+ as the foundation model and 0.8031 using ShenNong. This highlights the superior capability of TCM-FTP in modeling the intricate relationships between symptoms and herbs in prescriptions. Unlike previous approaches, TCM-FTP also includes herb dosage prediction, which is crucial in TCM due to the significant impact of dosage combinations. As shown in Table II, TCM-FTP achieves a much lower NMSE compared to the baseline using average statistics (0.0604 for TCM-FTP versus 0.0754 for Mengzi (T5-base) [26]). This enhances the practicality of TCM-FTP.
Compared to the pre-trained language model Mengzi (T5-base) [26], TCM-FTP significantly enhances prediction performance with improved precision (from 0.7332 to 0.7951), recall (from 0.7474 to 0.8113), and F1-score (from 0.7403 to 0.8031). Performances on GPT-3.5 and GPT-4.0 [16] were notably poor, indicating the limitations of general LLMs in tasks requiring specialized knowledge. In comparison to other TCM prescription prediction models, TCM-FTP outperforms PTM [3], which leads among these baselines, followed by TCMPR [24] in accuracy, with KDHR [25] and PresRecST [10] showing poorer performance. PTM, a topic model-based approach, is computationally complex, while TCMPR, KDHR, and PresRecST are graph-based models sensitive to graph-related factors. TCM-FTP excels these approaches in herb recommendations due to advanced language modeling capability.
IV-B Expert Qualitative Evaluation
Since conducting expert qualitative evaluation (EQE) is time-consuming and labor-intensive for doctors, this study randomly selected 20 data points from the test set for evaluation. As shown in Table III, the TCM-FTP model outperformed other baseline models across all metrics, achieving an average SHE score of 4.08 (standard deviation 0.5628) and an average HC score of 4.03 (standard deviation 0.5404). This indicates that TCM-FTP can generate effective prescriptions that adhere to TCM compatibility principles. In contrast, GPT-3.5 and GPT-4.0 [16] scored below 3 for both SHE and HC, with total scores of 5.75 and 5.89, respectively, significantly lower than TCM-FTP and TCMPR [24]. This demonstrates that GPT-3.5 and GPT-4.0 perform poorly in generating prescriptions that adhere to TCM principles. TCMPR performed relatively well, with SHE and HC scores of 3.56 (standard deviation 0.6715) and 3.54 (standard deviation 0.6578), respectively, but still fell short of TCM-FTP. Overall, TCM-FTP excelled in prescription generation, receiving higher approval from doctors and showing less score variability, indicating better stability and robustness.
TABLE III: Comparative Results of expert evaluation. The table records the mean scores (standard deviation).
Models GPT-3.5 GPT-4.0 [16] TCMPR [24] TCM-FTP(ours)
SHE 2.91 (0.7926) 2.93 (0.8319) 3.56 (0.6715) 4.08 (0.5628)
HC 2.84 (0.9819) 2.96 (0.9203) 3.54 (0.6578) 4.03 (0.5404)
Total 5.75 5.89 7.1 8.11
Figure 3: Two Specific Case Analyses. We present two specific test cases, including inputs, outputs, and predictions of the models with evaluations of experts. Herb names/dosage weights are marked in red for corrected ones.
IV-C Case Study
In order to visually showcase the predictive performance and capabilities of our TCM-FTP, we obtain the predicted herb results formed by each model on the test set and had them evaluated by TCM experts. Fig. 3 illustrates the results for two cases, including the input (chief complaints, present medical history), output (actual prescriptions provided by the doctor), and the predicted prescription results from TCM-FTP (ShenNong) and Mengzi (T5 base) [26] models, as well as GPT-3.5 and GPT-4.0 [16] predictions. Herb names/dosage weights marked in red indicate that the model’s prediction match the actual prescription provided by the doctor. Additionally, professional evaluations of each model are presented, ranking them using the “>” symbol (with models ranked higher indicating better performance). Results show that TCM-FTP (ShenNong) achieves strong prediction accuracy, closely matching doctors’ herb labels and dosages after fine-tuning, with clinically relevant dosage alignment. In contrast, GPT-3.5 and GPT-4.0 showed poor predictions due to a lack of specialized TCM training. This underscores the value of fine-tuning LLMs on high-quality, domain-specific data for optimal performance in specialized fields.
V CONCLUSION
To deal with the lack of high-quality datasets and to improve the performance in TCM prescription predictions, we build a TCM prescription dataset DigestDS and propose TCM-FTP to fine-tune LLMs to predict herbs with the corresponding dosages. The experimental results reveal the remarkable effectiveness of TCM-FTP, surpassing previous methods by large margins in precision, recall, and F1-score. Additionally, our method achieved the best results in NMSE, effectively forming accurate herb and dosage predictions. In future work, we will continue to incorporate domain knowledge into model construction to further enhance performance, aiming to develop a practically usable prescription prediction model.
VI ACKNOWLEDGEMENT
This work is partially supported by the Natural Science Foundation of Beijing (No. L232033), the National Natural Science Foundation of China (Nos. 82374302 and U23B2062), the Fundamental Research Funds for the Central Universities (No. 2024YJS032) and Hong Kong Research Grants Council under grant (No. 16204920). We gratefully acknowledge the support of NVIDIA Corporation with the GPU computational resources used for this research.